1. How much will you have in 10 years if you invest $20,000 at 6% compounded
annually?
2. How much must you invest today at 6% compounded annually so that you will
have $20,000 in 10 years?
3. How much will you have in 10 years if you invest $20,000 at 6% compounded
quarterly?
4. How much must you invest today at 6% compounded quarterly so that you will
have $20,000 in 10 years?
5. How much must you invest today at 6% compounded monthly so that you will
have $20,000 in 10 years?
6. How much will you have in 10 years if you invest $20,000 today at 6%
compounded monthly?

Found 2 solutions by ikleyn, Boreal:
Answer by ikleyn(52784)

.

One and ONLY ONE problem/question per post


It is the RULE and the POLICY of this forum.



Answer by Boreal(15235)


20000(1+0.06)^10=$35,816.95
the other way is x=20000/(1.06^10)=$11,167.90
the third is 20000*(1.015)^40=$36,280.37. divide inside interest by 4 and multiply outside exponent by 4 since this is 4 times a year
next one is x* (1+(0.06/4))^40=20000;
x=$11,025.25. You don't need to invest as much since more compounding times is more interest

the fifth is x*(1+.06/12)^120 =20000, and this will be $10,992.65
going the other way, one will have 20000*(1+(.06/12))^120, or $36,387.93